Circular No. 9004

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V5580 SAGITTARII = NOVA SAGITTARII 2008 No. 2 W. Liller, Viña del Mar, Chile, reports his discovery of an apparent

nova (mag ≈ 8.0) on two photographs taken with Tech Pan film and an orange filter on Nov. 29.040 UT; the variable was also present on Nov. 23.037 at magnitude ≈ 10.3, but was not visible (mag > 11.0) on Nov. 20.035. The new object is located at α = 18^{h}22^{m}.0, δ = −28^{o}03′ (equinox 2000.0). Liller adds that a pair of objective-grating spectrograms (yielding 44.0 nm/mm at Hα) taken on Nov. 30.013 clearly shows that the variable has a very strong Hα emission with a FWHM of ≈ 2500 km/s, quite typical of a classical nova several days after peak brightness. Because of the low altitude (∼ 17^{o}) of the apparent nova at the time of observation, the extinction in the blue-green portion of the spectra amounted to an estimated two magnitudes; for this reason, as well as to eliminate overlapping orders, an orange filter was used (with Tech Pan film). No other lines were visible. A direct photograph (orange filter, Tech Pan film) taken at Nov. 29.999 shows that the apparent nova’s approximate magnitude was 7.8, virtually unchanged from that of the discovery image. Note that the new object is located 1′.4 south-southeast of the tenth-magnitude star GSC 6852-3160. Liller reports mag 8.3 ± 0.1 for the nova from CCD images (with a red filter) obtained on Dec. 5.044 and 6.038; a better position measured by Liller is α = 18^{h}22^{m}01^{s}.39 ± 0^{s}.05, δ = −28^{o}02′39′′.8 ± 0′′.7.

G. W. Christie and T. Natusch, Auckland Observatory, report the following precise position for the nova (an average from three images taken with a 0.4-m Schmidt-Cassegrain reflector during Dec. 7.36, with the nova at altitude < 10^{o}, its magnitude being R = 9.5–9.9): α = 18^{h}22^{m}01^{s}.50 ± 0^{s}.03, δ = −28^{o}02′39′′.6 ± 0′′.5.

N. Samus and E. V. Kazarovets report that the GCVS team assigns the designation V5580 Sgr to this nova.

TIME ADJUSTMENT ON 2008 DECEMBER 31 Bulletin C36 of the International Earth Rotation and Reference Sys-

tems Service announces that a positive leap second will be introduced such that the sequence of UTC second markers will be: 2008 Dec. 31^{d}23^{h}59^{m}59^{s}, 31^{d}23^{h}59^{m}60^{s}, 2009 Jan. 1^{d}00^{h}00^{m}00^{s}. Beginning 2009 Jan. 1, the difference UTC–TAI = −34 s (UTC–TAI has been −33 s since 2006 Jan. 1; cf. IAUC 8567).

2008 December 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green